Speed of an Airplane. Two planes are 6,000 miles apart,and their speeds differ by 200 mph. They travel toward each other and meet in 5 hours. Find the speed of the slower plane.
Answer by mananth(16946)

Plane I x km/h
Plane II x+ 200 km/h
..
They are moving towards each other

So add up their speed.

combined speed = x+ 1 x + 200
( 2 x + 200 )
Time = 5 hours
Distance = 6000 miles
Distance = speed * time
( 2 x 200 )* 5 = 6000
10.00 x + 1000 = 6000
10 x = 6000 + -1000
10 x = 5000
/ 10
x= 500 mph
Plane I = 500 mph
Plane II 500 + 200 = 700 mph
m.ananth@hotmail.ca